Hello I need help anyone can someone explain n give me the answers


In the following geometric sequence,find the missing terms
3,?,?,15
Write the missing terms of the geometric sequence
3,_,_,15


(Type the exact answer,using radicals as needed)

Answer by Boreal(15235)

a1=3
a4=15
a4=3*r^3---------3r^(n-1) where n=4
3r^3=15
r^3=5
r= 5 ^(1/3)
a2=3*5^(1/3)
a3=3*5^(2/3)